### 08 Jan 2024 v2

# Scalable architecture for trapped-ion quantum computing using RF traps and dynamic optical potentials

Abstract
Qubits based on ions trapped in linear radio-frequency traps form a successful platform for quantum computing, due to their high fidelity of operations, all-to-all connectivity and degree of local control. In principle there is no fundamental limit to the number of ion-based qubits that can be confined in a single 1D register. However, in practice there are two main issues associated with long trapped-ion crystals, that stem from the ‘softening’ of their modes of motion, upon scaling up: high heating rates of the ions’ motion, and a dense motional spectrum; both impede the performance of high-fidelity qubit operations. Here we propose a holistic, scalable architecture for quantum computing with large ion-crystals that overcomes these issues. Our method relies on dynamically-operated optical potentials, that instantaneously segment the ion-crystal into cells of a manageable size. We show that these cells behave as nearly independent quantum registers, allowing for parallel entangling gates on all cells. The ability to reconfigure the optical potentials guarantees connectivity across the full ion-crystal, and also enables efficient mid-circuit measurements. We study the implementation of large-scale parallel multi-qubit entangling gates that operate simultaneously on all cells, and present a protocol to compensate for crosstalk errors, enabling full-scale usage of an extensively large register. We illustrate that this architecture is advantageous both for fault-tolerant digital quantum computation and for analog quantum simulations.
I Introduction
Trapped ions have ideal properties to be used as qubits for quantum computing (QC); they feature long coherence times, efficient state preparation and detection techniques, and a high degree of connectivity [1, 2, 3, 4, 5]. A quantum register of 1000s of qubits, or more, can be formed, for example, by utilizing an equally spaced crystal of ions in a linear RF Paul trap. Indeed recent years have seen many experimental attempts to work with increasingly larger trapped-ion registers [6, 7, 8, 9].
However, there are two practical issues associated with large ion-crystals that impede progress in this direction. The first is heating rates; as the number of ions in the crystal, $N$, increases, heating of the ions’ motional modes due to electric field noise drastically increases. Of particular concern is the axial center-of-mass (COM) mode, whose frequency typically decreases as $1/N$. This mode is especially vulnerable, as electric field noise tends to be spatially uniform and to target low-frequency modes [10, 11]. The resulting significant heating rates prohibit the implementation of high fidelity qubit operations, and might destabilize the ion-crystal.
The second issue with large ion-crystals is spectral crowding. As the size of the crystal increases, the frequency of adjacent motional modes becomes tightly spaced. For large ion-crystals with a dense mode spectrum, resolving individual modes becomes challenging; this complicates the implementation of entangling gates between two or more ions, which involves exciting their common motion. Granted, spectral control methods [12, 13, 14] allow for simultaneously targeting multiple modes to achieve desired qubit couplings with high fidelity. While such methods are promising for a moderate number of ions (up to 100s), they do not generally provide a scalable solution for arbitrarily large ion-crystals. First, the optimal control problems that must be solved to implement these methods becomes intractable for large $N$. Moreover, there is strong evidence that the minimum achievable gate time is set by the smallest frequency spacing among the motional modes [14]. This implies that gate times scale at least as $N^{2}$, making large ion-crystals prohibitively slow for quantum computations.
One active direction for scaling ion traps is the quantum charge-coupled device (QCCD) architecture [15, 16]. This setup involves many spatially separated trapping sites, each containing a small number of ions, where communication between sites is done by shuttling individual ions. Inevitably, this comes at the expense of high overhead in hardware and long circuit duration dominated by ion-shuttling and ion-cooling times. An alternative scale-up approach is using photonic-interconnects in order to link small-scale ion-crystals [17, 18, 19]. This method likewise involves a high overhead, due to the currently low entanglement rate via the interconnect, leading to slow operations.
Here we propose a scalable architecture for QC based on trapped-ion qubits that maintains the advantages of a long ion-crystal while circumventing its challenges. In our proposal, an arbitrarily long ion-crystal is segmented into cells by means of dynamically operated optical potentials, e.g. optical tweezers. In this way, the crystal’s motional mode structure is modified such that heating rates reflect only the cell size, and not $N$. Furthermore, programmable high-fidelity multi-qubit entangling gates can be implemented independently within each cell simultaneously. In addition, we show that mid-circuit measurements, a cornerstone of quantum error correction (QEC) techniques, as well as other central quantum computational tools, are straightforward to implement in an optically segmented ion-crystal.
In this work we generalize the method presented in Ref. [14], and employ qubit-local driving fields that implement programmable multi-qubit entangling gates. The method in Ref. [14] provides a clear physical intuition, an efficient procedure for designing gate drive parameters, and a thorough scaling analysis - all of which are relevant to this proposal.
The integration of optical tweezers into ion traps is an active area of research [20, 21, 22, 23, 24, 25]. Specifically Refs. [22, 23, 24, 25] make use of optical tweezers in order to generate target entanglement operations. In Ref. [20] the authors use tweezers for parallel two-qubit entangling gates, the system’s scalability is discussed however there is no treatment of large ion segments or ion-crystal heating rates. In Ref. [21] the authors consider segmenting an ion-crystal using tweezers, however do not analyze generating large-scale entanglement or connectivity between the segments. Mid-circuit measurements are not discussed in any of the above works. Unlike previous studies, here we detail how to achieve a holistic architecture, incorporating parallel and large-scale entangling operations and mid-circuit measurements, that are needed for large scale QC.
The remainder of this paper is organized as follows. In section II we present an overview of the proposed architecture. In section III we sketch two examples of applications of quantum information processing tasks that are amenable to our architecture: quantum simulation and quantum error correction. An in-depth analysis of our proposed architecture follows. Namely, in section IV we derive the spectral properties of optically segmented traps. In section V we analyze the implication of optical segmentation on the ion-crystal’s heating rate. In section VI we present our method for designing high-fidelity multi-qubit logical operations that are not hindered by unwanted crosstalk. Lastly, in section VII we present a protocol for performing mid-circuit measurements.
II Proposed architecture
There are well-established techniques for preparation, control, and measurement of trapped-ion based qubits for small-scale registers [26, 27]. We therefore focus mostly on unique aspects of our architecture. We propose a QC architecture in which a long trapped-ion crystal, made of $N$ ions, is segmented into $S$ cells, with each cell containing $C$ computational qubits available for quantum computation and simulation, generating a $C\cdot S$ qubit register. The segments are formed by placing $B_{A}$ barrier ions between adjacent cells, that are illuminated by an optical trapping potential, i.e. optical tweezers, made of tightly focused laser beams [28, 29], that provide an additional local confinement for the barrier ions.
The key point enabling scalability is that due to the optical segmentation the cells behave as nearly independent quantum registers. Indeed, we show below that both challenges of scaling-up trapped-ion based QCs, i.e., ion-crystal stability and the performance of logical operations, both scale with the cell size, $C$, and not the total number cells, $S$. Thus the dynamically reconfigurable segmentation removes the fundamental limits on the number of cells and on the number of qubits in the ion-crystal, and accordingly on quantum circuit and simulation size.
Figure 1: Architecture for scalable trapped-ion quantum computing using RF traps and dynamic optical potentials. The ion-crystal (vertical direction) is segmented to cells by dynamically applying optical trapping potentials. Sequential applications of parallel quantum operations and reconfiguration of the optical trapping implements a large-scale quantum circuit on the entire ion-crystal. We highlight a section of the quantum circuit and the sequence used. As shown, optical trapping generates $S$ segments, each containing $C$ computational qubits (purple circles) that are separated by $B_{A}$ consecutive optically confined ‘barrier ions (purple full circles). Additional $B_{B}$ ions in each cell (red circles) are used to support additional segmentation configurations of the ion-crystal, and mid-circuit measurements that enable classical feedback (red lines) for the implementation of e.g. quantum error correction. The horizontal direction shows a typical mode of operation of the QC, namely logical operations, involving multi-qubit and single-qubit logical gates, $U_{s,n}^{\left(A\right)}$, are performed in each cell, $s$, at the $n$th computational step. The optical potentials are then switched to a different segmentation configuration, allowing for large-scale connectivity between cells of the previous configuration, $U_{s,n}^{\left(B\right)}$. An intermediate segmentation configuration, in which all $B_{A}$ and $B_{B}$ ions are optically confined, is used to accommodate for mid-circuit measurement, cooling of the ion-crystal and state re-initialization. Additional computational steps ($n+1$, etc.) can be performed as required, thus ultimately guarantying connectivity between all computational qubits.
A general overview of our scalable architecture, is presented schematically in Fig. 1. In favor of a simple presentation we focus on a section of the entire ion-crystal, and a section of the operations used. We explicitly highlight $C=8$ computational qubits (purple circles) per cell. In practice 10s of qubits may be placed in each cell, and the number of cells is theoretically unlimited (vertical dots), forming a large ion-crystal. The cells are separated by $B_{A}$ barrier ions (purple full circles), with the schematic showing $B_{A}=2$. $B_{B}$ additional ions (red circles) are placed in the center of each cell and are used for segment re-configuration and mid-circuit measurements, with the schematic showing $B_{B}=2$. Thus, the ion-crystal contains $N=\left(C+B_{A}+B_{B}\right)S+B_{A}$ trapped ions.
The figure shows a general mode of operation of the QC, with parallel operations along the vertical axis and sequential operations, ordered from left to right, that form computational steps. Specifically, barrier ions are strongly confined with optical trapping potentials (purple full circles), generating the $S=4$ segments in this example (configuration ‘A’). At the $n$th computational step, local and global driving fields simultaneously implement unitary operations, $U_{s,n}^{(A)}$, with $s=1,2,...,S$. These unitary operators implement components of the overall quantum algorithm, and may involve all $C$ computational qubits (purple circles) and $B_{B}$ auxiliary ions (red circles) which are not optically confined in this configuration.
Next all $B_{A}$ and $B_{B}$ barrier ions are optically confined, enabling mid-circuit measurement of the auxiliary ions as well as recooling of the ion-crystal and preparation of ions to be used as auxiliary ions (detailed below). Classical feedback (red lines), based on the mid-circuit measurement results, may be implemented at this step in order to influence the next layer of unitary evolution.
Next, the segmentation is reconfigured by strongly confining previously unconfined $B_{B}$ barrier ions and removing the confinement of the $B_{A}$ barrier ions, thus interchanging their roles. This dynamically and quickly generates a different configuration of the segments (configuration ‘B’). Unitary operators, $U_{s,n}^{(B)}$, can now be implemented and connect previously uncoupled qubits. Additional measurement and reconfiguration steps can be applied as required. We remark that here we have shown only two basic ion-crystal configurations (A and B), and have further considered here and below $B_{A}=B_{B}$, however various additional segmentation configurations, with different parameters (e.g. segment and barrier sizes), can be flexibly generated provided relevant ions are allocated as barrier ions.
In addition to the configurable optical trapping potentials, we make use of local independently-applied fields for driving programmable multi-qubit gates [14] ($U^{(A)}$s and $U^{(B)}$s in Fig. 1), as well as for mid-circuit measurements and state preparation.
III Application demonstrations
We sketch examples of applications that showcase the utility of our architecture. These examples rely on various features of our proposal, e.g. implementing large-scale multi-qubit entangling operations, register reconfiguration, and performing mid-circuit measurements. The methods underlying these features are thoroughly discussed and analyzed in the sections below.
We start by considering analog quantum simulations on our system, specifically a three dimensional (3D) quantum-spin model. In general, quantum simulations are considered a well-suited task for noisy intermediate-scale quantum (NISQ) era QCs [30], as simulation of quantum systems is a notoriously challenging task for classical computers, while quantum computers are considered naturally suited for it [31]. Indeed, numerous quantum systems have recently implemented impressive demonstrations of quantum simulations [32, 33, 34].
Quantum simulations using trapped-ion based qubits, have also been recently demonstrated [35, 36, 37, 38, 39, 40, 41, 42, 43], many of which take advantage of the unique all-to-all coupling present in these systems. In these works Ising-type interactions, that are inherent to trapped-ion systems, can be straightforwardly iterated and combined with single-qubit rotations in order to generate arbitrary XYZ-type spin-models [44, 45].
Our architecture is well-suited for quantum simulations, as our ability to design programmable multi-qubit entanglement manifests as the dimensionality and geometry of the simulated model [46], and the optical segmentation and reconfiguration provide a straightforward approach to simulating 3D systems. Such 3D systems are challenging to simulate on linear or planar quantum processors with short-range interactions, due to a large gate overhead required for embedding the 3D geometry.
Figure 2: Quantum simulation of a 3D rectangular Ising spin-model on an optically segmented ion-crystal. Computational qubits (purple circles) are coupled in cells (purple brackets) such that each cell holds two planes of the rectangular spin-model, and implements inter-plane (dark-blue and bright-blue links) and intra-plane (red links) couplings (a) Couplings of a single cell, holding two planes of the model, with $d=4$, used for both segmentation configurations. (b) Multiple cells form the 3D Ising model. The two optical configurations (‘A’, ‘B’ and purple brackets) enable interlaced coupling of the planes (solid red and dashed red, respectively), such that two entanglement operations implement the model’s Hamiltonian. (c) We design the control pulses that implement the nearest-neighbor couplings required for the Ising model, here for $d=4$ and $S=5$. The resulting coupling between all qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that implements the links shown in (a), for each cell. (d) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The implementation’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
For simplicity, we focus here on a 3D rectangular lattice of spins, though many other geometries can be realized, e.g. hexagonal. Figure 2 shows an implementation of such a $d\times d\times 2S$ rectangular Ising model. Specifically, Fig. 2(a) shows a single cell of the ion-crystal, that contains $2d^{2}$ computational qubits (purple circles), and encodes a pair of two-dimensional planes of the simulated system. Programmable entanglement gates implement the required couplings that generate the model’s geometry, i.e. inter-plane couplings (dark-blue and bright-blue) and intra-plane couplings (red). The model is made three dimensional, as shown in Fig. 2(b), by interlacing the vertical couplings between adjacent layers (solid red and dashed red), using the two segmentation configurations (purple brackets).
As an example, we consider a specific realization of our architecture, namely qubits implemented on the $4S_{\frac{1}{2}}$ Zeeman ground-state manifold of trapped ${}^{40}\text{Ca}^{+}$ ions, coupled by Raman transitions mediated by the $4P$ manifolds. We utilize this realization for a $d=4$ and $S=5$ nearest-neighbor Ising model (see further system details in the SM [47]). We design the control pulses that drive the computational qubits, and generate the required entanglement operations for simultaneous $XX$-coupling of the qubits, in the entire ion-crystal, similarly to the couplings shown in Fig. 2(a). For simplicity, we use a uniform, unity coupling between all nearest-neighbor qubits. Our design protocol is presented and discussed below.
We calculate the coupling terms that result from the control pulses that we designed, and plot them in Fig. 2(c), with each point on the plot representing a qubit-qubit coupling between qubits of the corresponding indices (horizontal and vertical axes). The block-diagonal structure that is seen reflects the underlying segmented structure of the ion-crystal. The relative difference between the ideal and the resulting qubit-qubit couplings is shown in Fig. 2(d) in log-scale exhibiting a low coupling error that indicates a high-fidelity and accurate simulation. Indeed, the overall performance of our implementation, simultaneously coupling $160$ computational qubits, is evaluated with an infidelity that is better (lower) than $10^{-4}$ per cell.
We note that with this encoding, a time-step of an $XX$ nearest-neighbor 3D rectangular Ising model can be implemented with only two sequential multi-qubit entanglement gates, such as that shown in Fig. 2(c), which, for $d=4$ and $S=5$, would otherwise require the application of 384 sequential two-qubit gates between arbitrary pairs in the quantum register (in general $\mathcal{O}\left(d^{2}S\right)$ two-qubit gates). By considering more elaborate models, e.g. next-nearest neighbor interactions within the two-dimensional planes, non-uniform couplings between the planes, or larger systems, the two-qubit gate count will increase, while with our method two multi-qubit gates still suffice.
The second example showcases a path towards fault-tolerant quantum computation [48, 49, 50]. As shown in 3(a), each segment (purple brackets, ‘A’) is utilized to encode a single logical error-corrected qubit using a distance-five surface code. The code uses 25 computational qubits to store the logical state, and 12 auxiliary (ancilla) qubits to readout the values of both $X$ and $Z$ stabilizer measurements. The entire set of $X$ stabilizer measurements can be implemented in one step, using a single multi-qubit entanglement gate and mid-circuit measurements. Then, the auxiliary qubits are reinitialized, after which the entire set of $Z$ stabilizers is similarly performed. Clearly cell segmentation does not require 12 barrier ions, thus here most of the 12 auxiliary qubits are implemented by computational qubits in the cell.
Figure 3: Quantum error correction code on an optically segmented ion-crystal. (a) We implement several distance-five logical qubits, shown pictorially as grids of $5\times 5$ qubits (purple circles). Each cell in configuration ‘A’ (purple brackets) houses a single logical qubit (labeled 0 - 12 and 25 - 36) and twelve auxiliary qubits (labeled 13 - 24). Programmable multi-qubit gates generate the required entanglement structure for stabilizer measurements on each logical qubit in parallel. Here we show an implementation that generates $X$-stabilizers (light-green plaquettes and blue lines). $Z$-stabilizers may be implemented with an additional gate (light-red plaquettes). Connectivity across the entire register is enabled by configuration ‘B’ (light-purple brackets) such that each cell in this configuration holds data qubits of two adjacent logical qubits. This enables logical entangling operations between any neighboring encoded qubits via stabilizer measurements formed on the border between the two surface codes (red plaqeuttes and blue lines), using the auxiliary ions in this configuration (labeled 17’-19’). (b) Example of programmable multi-qubit gate generating $X$ and $Z$ stabilizers on five logical qubits, implemented on a segmented ion-crystal with $S=5$, $C=37$ and $B=3$. We design the control pulses that implement the entanglement gates required for the various stabilizers. The resulting qubit-qubit couplings between corresponding qubit pairs (horizontal and vertical axes) is shown (color), exhibiting a block-diagonal structure that reflects the formation of five logical qubits. (c) Deviation of the qubit-qubit coupling from the ideal structure (log scale), showing a low error, that is not limited by crosstalk between adjacent cells, and enables scaling up. The gate’s infidelity scales as the deviations squared, and is here evaluated to $10^{-4}$ per cell.
Naturally, these stabilizer measurements can be performed across all logical qubits in parallel. Note that not all segments must implement the same operation at any given time; some may, for example, implement $X$ stabilizer measurements while others implement $Z$ stabilizer measurements. Figure 3(b) shows a specific example involving five logical qubits, and multi-qubit entanglement gates, designed to perform simultaneous $X$ and $Z$ stabilizer couplings, in different cells, throughout the entire ion-crystal. The relative difference between the ideal and the resulting qubit-qubit couplings is shown in Fig. 3(c) in log-scale, exhibiting a low error that indicates a low infidelity that is better (lower) than $10^{-4}$ per cell.
Entanglement between two logical qubits, implemented on independent surface codes, can be achieved, for example, using ‘lattice surgery’ [51, 52]. In this method, stabilizers formed along the border between two surface codes are measured, projecting the joint state of the logical qubits to an entangled basis. In our case, this is naturally enabled by re-configuring the optical segmentation (light-purple brackets, ‘B’) such that ions from neighboring cells can interact, as shown schematically in Fig. 3 (blue lines and red plaquettes in the center of the ’B’ configuration). Following this operation, the segmentation reverts to its original configuration, and additional stabilizers are measured. The overall result of this procedure is an entangling logical $XX$ operation between the adjacent surface codes.
We emphasize that even though five cells are used in this example, our architecture’s scalability enables using considerably more cells in order to implement many logical qubits. This example shows one potential route to fault tolerance using our architecture; however many different variants could be considered. First the distance of the code, and thus the number of physical qubits required, should be determined by the demands on the overall logical error rate. Moreover, the choice of the surface code itself, while motivated by its high fault tolerance threshold, is not the only possible choice. In fact, there are other high-threshold codes, with potentially better encoding rates, that may be suitable as well [53]. Constructing an optimal protocol for fault tolerant QC within our architecture, including a prescription for performing logical non-Clifford operations, is a subject for further study.
IV Spectral properties of optically-segmented ion-crystals
In general, similar to alternative scale-up methods, segmentation decouples cells from each other, reshaping the well-known all-to-all coupling of trapped-ion crystals to local all-to-all couplings within cells of manageable sizes of tens of ions. As we show below, this segmentation preserves the ion-crystal’s stability and enables programmable multi-qubit quantum gates that act simultaneously and independently within the different cells. Another unique feature of our scale-up strategy, compared to other segmentation approaches, stems from the ability to dynamically and quickly reconfigure the applied segmentation, generating new cells that combine and couple previously decoupled qubits, thus enabling fast and large-scale connectivity within the whole ion-crystal.
We analyze the effect of optical segmentation on the ion-crystal, specifically on collective modes of motion of the ion-crystal in the axial and transverse directions. To this end we assume that the barrier ions are illuminated by beams that generate an optical trapping potential (o.t.p), inducing the trapping frequency [54],
$\omega_{\text{o.t.p}}=\sqrt{\frac{2\text{Re}[\alpha\left(\lambda\right)]}{m}}|E|,$ (1)
where $\alpha\left(\lambda\right)$ is the wavelength-dependent polarizibility of an illuminated ion, $m$ is its mass, and $E$ is the field strength. The impact of this potential is best quantified by comparing it to another important parameter in the system, namely the characteristic frequency-scale associated with the Coulomb interaction between adjacent ions [20],
$\nu=\sqrt{\frac{e^{2}}{4\pi\epsilon_{0}md^{3}}},$ (2)
with $e$ the elementary electron charge, $\epsilon_{0}$ the vacuum permittivity and $d$ is the inter-ion distance of an equidistant ion-crystal. As we show below, a strong optical potential, $\omega_{\text{o.t.p}}>\nu$, drastically changes the motional mode structure of the chain. By incorporating the optical trapping potential with common methods to analyze ion-crystal motion [55], we obtain the normal-mode frequencies and structure in the axial and transverse directions (for further details see the SM [47]).
We first consider the effect of the optical trapping on the axial modes of motion of the ion-crystal, that lie at the lower parts of the motional spectrum and are typically more prone to heating. When considering the joint motional mode structure of two identical ion-crystals, in two completely separate traps, we observe a degenerate mode structure, with each motional frequency appearing twice, and each motional mode being localized at one of the ion-crystals. Thus we intuitively expect that the motional spectrum of a well-decoupled optically segmented ion-crystal will form approximately degenerate bands.
Figure 4 demonstrates this by showing the axial motional spectrum of a $N=231$ long ion-crystal for a varying optical trapping potential that segments the ion-crystal into $S=6$ cells, separated by $B_{A}=3$ barrier ions, such that each cell contains 35 ions (out of which $C=32$ are computational qubits and $B_{B}=3$ acts as barrier ions for segmentation configuration ‘B’). As the optical trapping potential increases the axial spectrum (blue) forms 35 bands of $S$ modes each.
Figure 4: Axial spectrum of an ion-crystal with N=231 ions, segmented into $S=6$ cells. Cells are separated by $B_{A}=3$ barrier ions and include $C+B_{B}=35$ ions each. The axial frequencies (blue) are shown for various trapping optical potentials, $\omega_{\text{o.t.p}}$ and are normalized by the characteristic Coulomb scale, $\nu$. For comparison, an additional axial spectrum of an independent ion-crystal with 35 ions is shown (dashed orange). For $\omega_{\text{o.t.p}}>\nu$ the axial spectrum of the optically segmented trap forms bands which are located at the frequencies of the independent cell. As $\omega_{\text{o.t.p}}$ increases, high-frequency bands form, which are due to the $B_{A}$ barrier ions. We highlight, $\omega_{\text{o.t.p}}/\nu\approx 2.1$ (vertical line), used throughout the text. The inset is a zoom-in on the center-of-mass band (vertical axis in log-scale), showing that at $\omega_{\text{o.t.p}}=2\nu$ the segmented crystal’s modes resemble the independent cell, implying that heating is dictated by the properties of a single cell rather than the whole, multi-cell, crystal.
Indeed the formation of bands heralds the decoupling of the motion of cells from each other, such that band frequencies resemble those of an independent, un-segmented, ion-crystal containing $C+B_{B}$ ions (orange), with each mode in the band being a superposition of local excitations of the corresponding mode of the independent cell. Thus, the width of each band marks the relevant rate in which a motional excitation traverses from one cell to another. That is, an excitation of a local mode in a single cell is composed of a superposition of all modes of the band, which then disperse at a time scale that is inversely proportional to the band width.
Figure 4 further shows $B_{A}$ high-frequency bands, containing $S+1$ modes, associated with the motion of $\left(S+1\right)B_{A}$ barrier ions. Crucially these bands are separated from the ‘bulk’ ions such that their motion is essentially decoupled from the bulk. We exploit this fact below in order to implement mid-circuit measurements and cooling of the barrier ions.
Figure 5 shows the transverse (radial) band structure of the same $N=231$ ions system with $\omega_{\text{o.t.p}}\approx 2.1\nu$ (dashed vertical line in Fig. 4). Similarly to the axial direction, optical segmentation results in the formation of $C+B_{B}$ bulk-bands as well as additional $B_{A}$ bands associated with the motion of barrier ions (blue). The average frequencies of the bands resemble those of an independent cell (orange), also shown in the inset (bottom right) that presents a zoom-in on the last four computational qubit bands, showing narrow and well-separated bands. For transverse modes of motion the band index, $b$ (horizontal axis), ordered according to the band’s frequency, is opposite to the mode’s wave vector, i.e. large $b$’s represent long-wavelength modes of motion.
Transverse modes of motion are used in order to generate entanglement between computational qubits, by mediating qubit-qubit interaction via spin-dependent motion. Typically the entanglement operation duration is inversely related to the frequency difference between adjacent modes of motion [14].
Here, the frequency gap between bands, $\Delta\omega_{b}=\omega_{b}-\omega_{b-1}$, with $\omega_{b}$ the mean frequency of the modes of bulk bands, $b=1,...,C+B_{B}$, marks the typical interaction rate between ions within the same cell. Thus a simple estimation of unwanted crosstalk in the entanglement operations between cells is given by $\varepsilon_{\text{BW},b}=\frac{\text{BW}_{b}}{2\Delta\omega_{b}}$, with $\text{BW}_{b}$ the bandwidth of band $b$, that sets the coupling rate between cells. The resulting estimate is presented in the inset (top left) of Fig. 5 (orange), showing that high $b$ bands, associated with long-wavelength motion across cells, results in a higher crosstalk, at the few percent level. A more precise estimate of the crosstalk is also shown (blue line and green points) and discussed below.
Figure 5: Radial modes spectrum of an optically-segmented ion-crystal containing $N=231$ ions, and the resulting cell crosstalk. Due to the segmentation the mode frequencies (blue) form $C+B_{B}$ bands, which imitate the structure of an unsegmented independent cell (orange). Additional $B_{A}$ high-frequency bands associated with the barrier ions are formed as well. The bottom-right inset shows a zoom-in on the last four bulk bands. The bands are well separated and narrow in frequency. The top-left inset presents estimations of unwanted crosstalk between cells during simultaneous multi-qubit operations, mediated by the various bulk bands. These estimations are calculated by considering the system’s spectrum, $\varepsilon_{\text{BW},b}$ (orange) or mode-structure considerations, $\varepsilon_{J,b}$ making use of the modes of the entire ion-crystal (blue lines), or only neighboring segments (green stars). Both predict a $\sim 2.5\%$ crosstalk due to mainly long-wavelength motion of ions in the cells, associated with high bands. Mitigation of the crosstalk is presented in section IV in the main text.
Optically confined ions are not used as computational qubits since optical confinement significantly degrades their coherence due to spontaneous photon scattering and uncontrolled light-shifts. For example, we consider ${}^{40}\text{Ca}^{+}$ ions that are illuminated by a confining optical field at $400$ nm with a beam that has a diameter of 1 $1\text{ }\mu\text{m}$, off-resonant with the $S\leftrightarrow P$ transition. In this setup, a harmonic confinement frequency of $\omega_{\text{o.t.p}}\approx 2.1\nu=(2\pi)*1.75$ MHz, used throughout our analysis, requires an optical power of $20$ mw illuminating each barrier ion. The resulting spontaneous photon scattering rate is approximately $3$ kHz. Photons scattered from barrier ions are significantly detuned (by $10$’s of THz) with respect to the $S\leftrightarrow P$ transition in neighboring computational qubits, which thus acquire only a negligible decoherence rate due to this process.
V Heating rates of optically-segmented ion-crystals
A crucial criteria for successfully scaling-up ion-crystals is maintaining crystal stability. In the standard approach to trapped-ion crystals this challenge comes about as a heating rate of the crystal that scales unfavorably with $N$. Specifically, in un-segmented ion-crystals as $N$ increases, the frequencies of low-lying long-wavelength modes of motion decrease, and become more susceptible to electrical noise, leading to degradation of coherent operations, destabilization of the ion-crystal and its eventual melting. Utilizing optical trapping potentials generates cell-local heating rates, which only depend on the cell size, $C$, and not on the total number of ions $N$ (or $S$), potentially allowing for arbitrarily long ion-crystals.
This is shown graphically in Fig. 4, since for $\omega_{\text{o.t.p}}\geq 2\nu$ the low-lying COM band is lifted and its frequency converges to the COM mode of an independent cell (inset), implying cell-dependent (and not $N$-dependent) heating rates. The COM modes remain gapped regardless of $N$.
We directly quantify the phonon excitation rate of the optically-segmented architecture. For a given motional mode, the phonon excitation due to the presence of some electric field noise, $\delta E(t,r)$,can be evaluated by Fermi’s golden rule [11], yielding,
$\Gamma^{(k)}=\frac{e^{2}}{4m\hbar\omega_{k}}S_{E}^{(k)}(\omega_{k}),$ (3)
with $\omega_{k}$ the frequency of mode $k$, and $S_{E}^{(k)}(\omega_{k})$ is the spectral density of the electric field noise. The mode’s heating rate is $\hbar\omega_{k}\Gamma^{\left(k\right)}$.
The precise form of the spectral density function is a subject of much theoretical and experimental study [56, 57, 58, 59, 10, 11, 60], is system-dependent and its theoretical limits remains somewhat inconclusive. Yet it is generally agreed that it scales as, $S_{E}(\omega)\propto\omega^{-\alpha}D^{-\beta}T^{\gamma}$, with mode frequency $\omega$, ion-electrode distance $D$, and temperature, $T$, and $\alpha$, $\beta$ and $\gamma$ are scaling exponents. Motivated by many experimental results, here we assume that the noisy electric field is spatially uniform along the ion-crystal and that $\alpha=1$.
Using these consideration we obtain the system’s excitation rate (see details in the SM [47]),
$\Gamma=\sum_{k}\Gamma^{(k)}\propto\sum_{k}\frac{e^{2}}{4m\hbar\omega_{k}^{1+\alpha}}\sum_{i,j}A_{k}^{(i)}A_{k}^{(j)},$ (4)
with $\Gamma^{(k)}$ the excitation rate of the $k$th mode and $A$ an orthogonal axial mode-matrix, such that $A_{k}^{(i)}$ is the participation of the $i$th ion in the $k$th axial mode of motion, obtained by analyzing the axial mode structure [55, 47].
Due to the summation over $k$, our formulation of $\Gamma$ in Eq. (4) is extensive, e.g. for two identical and independent copies of an ion-crystal, $\Gamma$ will be evaluated as twice the value of a single copy of the ion-crystal. Thus ideal decoupling between cells in the optically segmented ion-crystal is expected to show-up as proportionality of $\Gamma$ to $S$.
Figure 6 shows the predicted excitation rate of the optically segmented ion-crystal, due to Eq. (4), for various values of the cell size, $C$ (horizontal axis), and number of segments, $S$ (colors). The barrier size is set to $B_{A}=3$ and each cell contains $C+B_{B}$ ions, with $B_{B}=3$. Excitation rates are calculated for ion-crystals reaching $N>1300$ ions. The figure shows, $\Gamma/\left(\Gamma_{\text{ind}}S\right)$ (vertical axis), such that the excitation rates are normalized by the number of segments, $S$, and by the excitation rate of an independent unsegmented ion-crystal containing $C+B_{B}$ ions, $\Gamma_{\text{ind}}$. We observe that the normalized excitation rates are all of order unity and converge towards 1 as $C$ increases, with a negligible dependence on the number of segments. Thus we conclude that the cell’s stabilization is dominated by cell properties, and is specifically independent from other cells in the ion-crystal.
The main contribution to $\Gamma$ is given by the lowest frequency band, which is a COM motional band. This is intuitive as COM modes have the lowest motional frequencies, as well as a high overlap with the uniform electric field noise considered here, as compared to other modes which have negligible contributions. This is seen in the dashed lines of Fig. 6 in which we calculate $\Gamma$ using exclusively this single band (while $\Gamma_{\text{ind}}$ is still calculated in full).
Figure 6: Axial excitation rate of optically segmented ion-crystals. The excitation rates (vertical axis), $\Gamma$, are shown for various values of cell sizes, $C$ (horizontal axis) and number of segments, $S$ (colors). We normalize $\Gamma$ by the excitation rate of an independent single cell, $\Gamma_{\text{ind}}$ and by the number of segments, $S$. All resulting normalized rates are of order unity, showing that the stability of the optically segmented ion-crystal is dictated by $C$, and is largely independent of $S$. We repeat this process but with only considering the lowest band of motional modes in the expression for $\Gamma$ (dashed), without varying $\Gamma_{\text{ind}}$. The resemblance between this result and the full expression of $\Gamma$ (solid) shows that the lowest-lying band is the dominant contributor to the heating. This result is intuitive as this band is a COM motional band, having a high-overlap with a uniform electric field.
VI Parallel multi-qubit logic and crosstalk mitigation
In trapped-ion based quantum computers, logical gates between qubits, i.e. entanglement, is typically facilitated by spin-dependent forces acting on normal-modes of motion of the ion-crystal [61, 62]. Thus adequate control of the normal-mode structure is crucial for scaling-up the trapped-ion quantum register size. Specifically, utilizing many modes of motion simultaneously enables programmable long-range interactions [12, 13, 14], which are significant for efficient implementation of various quantum computational tasks [63, 64, 65, 66, 53]. However utilizing multiple modes of motion becomes increasingly challenging in large ion-crystals due to the overwhelming complexity of the required qubit drive, which comes about as a NP-hard optimization problem [12], as well as an apparent slow-down of the feasible gate duration [14] which, in absence of segmentation, scales as $N^{2}$. Thus segmentation plays here a crucial role as well in generating independent large, yet manageable, computational qubit cells.
The heuristic estimation for crosstalk error, provided by $\varepsilon_{\text{BW,b}}$ above, can be improved by a detailed consideration of the transverse mode-structure. Specifically, based on the quadratic-form of qubit-qubit entanglement discussed thoroughly in [14], the coupling term between ions $c$ ($c^{\prime}$) in cells $s$ ($s^{\prime}$), mediated by band $b$, is estimated as
$J_{b}^{\left(c,s\right),(c^{\prime},s^{\prime})}=\sum_{m=1}^{S}R_{\left(b,m\right)}^{\left(c,s\right)}R_{\left(b,m\right)}^{(c^{\prime},s^{\prime})},$ (5)
with $R$ an orthogonal radial mode-matrix (similar to $A$ above) such that $R_{\left(b,m\right)}^{\left(c,s\right)}$ is the participation of ion $c$ of cell $s$ in mode $m$ of band $b$ (specific examples of $J_{b}^{\left(c,s\right),(c^{\prime},s^{\prime})}$ are shown in the SM [47]). Using $J_{b}$ above, we form a more detailed crosstalk estimation, namely,
$\varepsilon_{J,b}=\frac{\max\limits_{c\neq c^{\prime},s\neq s^{\prime}}\left|J_{b}^{\left(c,s\right),(c^{\prime},s^{\prime})}\right|}{\max\limits_{c\neq c^{\prime},s}\left|J_{b}^{\left(c,s\right),(c^{\prime},s)}\right|},$ (6)
with $s,s^{\prime}$ maximized over all $S$ cells in the segmented ion-crystal and $c,c^{\prime}$ maximized over all $C$ computational qubits in each of the cells. In essence the estimation in Eq. (6) compares the coupling between ions in the same cell ($s=s^{\prime}$) to coupling between ions in difference cells ($s\neq s^{\prime}$). This estimation is also shown in the inset (top-left) of Fig. 5 (blue), exhibiting a similar behavior to that of $\varepsilon_{\text{BW},b}$, i.e. $\sim 2.5\%$ crosstalk mainly due to long-wavelength motional excitations of the cells. We also evaluate separately the contribution of nearest-neighbor cells, $\varepsilon_{J,b}^{\text{nn}}$, which follows the same definition as $\varepsilon_{J,b}$ in Eq. (6) but with $s^{\prime}=s\pm 1$ in the numerator. This estimation is shown as well in the inset (top-left) of Fig. 5 (green stars), showing an almost exact agreement with $\varepsilon_{J,b}$. Thus we conclude that the $\sim 2.5\%$ crosstalk is mainly due to unwanted coupling between adjacent segmented cells. Further analysis shows that in general 2-3 barrier ions, each confined with an optical trapping frequency of $\omega_{\text{o.t.p}}\approx\omega_{\text{rad}}$ are sufficient for maintaining low crosstalk levels (see the SM [47]).
A few percent crosstalk level, reflected by both $\varepsilon_{\text{BW},b}$ and $\varepsilon_{J,b}$, marks an efficient decoupling between cells. Nevertheless, the remaining coupling is too high for high-fidelity quantum operations, and restricts performing simultaneous logical gate on distinct cells. However this crosstalk level is low enough such that it can be mitigated perturbatively in a scalable manner. Here we specifically rely on the large-scale fast (LSF) method, presented in Ref. [14].
In essence, LSF makes use of a multi-tone drive in order to generate programmable multi-qubit gates, which naively involves solving a quadratically constrained optimization problem of dimension $N^{2}$, and generates the unitary gate, $\exp\left(\sum_{n,m=1}^{N}\varphi_{n,m}\sigma_{x}^{\left(n\right)}\sigma_{x}^{\left(m\right)}\right)$, with $\sigma_{x}^{\left(n\right)}$ a Pauli-$x$ operator acting on the $n$th qubit and $\varphi_{n,m}$ a target multi-qubit coupling matrix. The hardness of the optimization problem is mitigated in LSF with a key realization; namely, that for each fixed system, any non-vanishing drive solution to a special ‘zero-phase’ version of the optimization problem ($\varphi^{\left(0\right)}_{n,m}=0$) can be efficiently converted to a high-fidelity solution of any general problem. Finding such a zero-phase solution (ZPS) is generally a far easier task than directly solving the general problem.
The operational approach of LSF is naturally adopted to our architecture, making the design of multi-qubit entanglement, simultaneously in all cells, fast and scalable. Specifically, we use LSF in order to generate ZPSs for an effective ‘typical’ cell of the bulk of the ion-crystal (see the SM [47]). Up to small deviations that are corrected below, this ZPS is adequate to all of the ion-crystal’s cells. Using LSF, different multi-qubit gates are designed for each cell by converting the same ZPS. This classical computation can be performed fast and in parallel. After performing these steps we obtain a specific drive spectrum, for all computational and auxiliary qubits in the ion-crystal, that generates the required entanglement per-cell at once, $\exp\left(\sum_{s=1}^{S}\sum_{n,m=1}^{C}\varphi_{n,m}^{\left(s\right)}\sigma_{x}^{\left(n,s\right)}\sigma_{x}^{\left(m,s\right)}\right)$, with $\sigma_{x}^{\left(n,s\right)}$ a Pauli-$x$ operator acting on the $n$th qubit in cell $s$ and $\varphi_{n,m}^{\left(s\right)}$ the desired multi-qubit target of cell $s$.
As discussed above, the unwanted coupling between cells incurs a $\sim 2.5\%$ crosstalk error, which is dominated by nearest-neighbor cells. Here this brings about a typical coupling of $\sim 0.025\varphi_{n,m}^{\left(s\right)}$ between ions in distinct cells. Therefore we make use of an additional mitigation layer on top of LSF. Crucially, the nearest-neighbor structure of the crosstalk enables our mitigation technique to scale favorably, in terms of classical computation resources, and does not impose constraints on the ion-crystal size.
Specifically, we perform an iterative optimization by linearizing the quadratic constraints around the current solution, with the initial solution given by LSF as discussed above. The resulting linear equations provide conditions that resolve crosstalk as well as inaccuracies of the LSF solution that originate from the assumption of a typical cell. We do so locally, i.e. the linear conditions are formed for only two adjacent cells at a time, generating $C\left(C-1\right)$ linear constraints that improve the LSF results and $3C^{2}$ linear constraints that correspond to crosstalk terms. The iterative optimization is stopped when a target infidelity is reached. Crucially since the linear equations are local, relating only four adjacent cells at a time, they involve only $\mathcal{O}\left(C^{2}\right)$ linear constraints, allowing parallel optimization of the next four adjacent cells [47].
An additional condition for a high-fidelity process is decoupling of the qubits from the modes of motion at the end of the entanglement operation. This is typically satisfied by forming constraints for each of the $N$ modes of motion independently. Here however such an approach will be a hurdle to scalability since it will be infeasible to satisfy a large $N$ number of constraints. Instead, we make use of only $C+B_{A,B}$ linear constraints of the typical cell, and supplement these with additional robustness properties that make the decoupling insensitive to the mode frequency inaccuracies, described in Ref. [67], such that each constraint enforces the decoupling of an entire band at once. The error in our gates is therefore expected to have a contribution from imperfect disentanglement of motion from qubits.
Figure 7: Performance of the crosstalk mitigation method as a function of optical trapping potential $\omega_{\text{o.t.p}}$, evaluated on the entanglement used to implement a simulation of a 3D rectangular Ising model. Top: Error per cell, using Eq. (7), as a proxy for the entanglement infidelity of the entire multi-qubit operation. For $\omega_{\text{o.t.p}}\gtrsim\omega_{\text{rad}}$ the crosstalk compensated solution (dark-blue) shows a low, $<10^{-4}$, infidelity. As expected, the initial LSF solution, based on the effective typical cell (orange) shows a $\sim 10^{-2}$ infidelity. The mitigation degrades in the weak optical confinement regime, due to long-range coupling between cells that is not considered in the crosstalk mitigation method. Indeed this degradation appears when considering long-range cells (dashed purple), and not seen when only considering nearest-neighbor and onsite coupling (dashed pink). Additional infidelity due to residual coupling to motional modes (dashed light-blue) exhibits a similar behavior. Bottom: The required total drive coupling, given in units of the lowest transverse mode frequency (bottom). The crosstalk compensated solution (dark-blue) only incurs a small, $\sim 10\%$, overhead on the initial LSF solution (orange).
We study the efficacy of our crosstalk mitigation method by considering a segmented ion-crystal system used for a 3D Ising model simulation, as that shown in Fig. 2 above, with $S=5$, $d=4$ and $B_{A}=B_{B}=3$. We design the control pulses that implement the required qubit-qubit couplings, for various optical trapping frequencies. The resulting operation infidelity is evaluated, to leading order in the qubit-qubit coupling deviations and in residual ion displacements, as,
$I=I_{\text{short}}+I_{\text{long}}+I_{\text{motion}}$ (7)
such that $I_{\text{short}}$ approximates the infidelity due an erroneous qubit-qubit coupling within each cell and crosstalk to nearest-neighbor cells, $I_{\text{long}}$ is the infidelity due to qubit-qubit coupling between cells that are not nearest-neighbors and $I_{\text{motion}}$ approximates the infidelity due to residual coupling between the qubits and the phonon modes of motion.
Figure 7(top) presents the infidelity, $I$, of the initial LSF solution, based on the typical cell (orange) and the crosstalk corrected solution (dark-blue). It exhibits an apparent cross-over between two regimes at $\omega_{\text{o.t.p}}=2\nu$, at which point the transverse mode band structure is manifested. At low optical confinement the performance of uncorrected, typical-cell solutions (orange) is low, such that our leading-order estimation breaks-down. At strong optical confinement this solution generates an infidelity of $\sim 10^{-2}$, as expected from our analysis above. However the crosstalk-corrected solution (dark-blue) successfully generates solutions that have an infidelity that is better (lower) than $10^{-4}$ for the entire multi-qubit operation, throughout the strong optical confinement regime. We account for the different contributions to the infidelity of the crosstalk-corrected solution, given by Eq. (7). We note that $I_{\text{short}}$ (dashed pink), has a negligible contribution, regardless of optical confinement, while $I_{\text{long}}$ (dashed purple) reaches a low values only at the strong confinement regime. This behavior is expected as our optimization actively corrects for on-site and nearest-neighbor terms, and ignores other terms. We note that $I_{\text{long}}$ reaches a minimum at $\omega_{\text{o.t.p}}=2\nu$, and is slightly increased at higher optical trapping frequencies. This behavior is accounted by an effective screening of barrier ions that occurs when barrier bands are close to resonance with the bulk bands (see the SM [47]). Nevertheless, our analysis implies that strong optical confinement successfully decouples next nearest-neighbor (and further) cells, removing the need to take the entire ion-crystal into account, and enabling scalability. Similarly, due to a well-formed band structure, the motional infidelity, $I_{\text{motion}}$ (dashed light-blue), becomes small at strong optical confinement, and in general dominates the overall infidelity.
Figure 7(bottom) shows the total drive coupling, i.e. Rabi frequency, required to drive the initial LSF solution (orange) and the crosstalk corrected solution (dark-blue), given in terms of the frequency of the lowest radial mode, $\nu_{1}$, as a characteristic scale. The compensated solution incurs a small overhead of $\sim 10\%$ on the initial solution.
We remark that elements of our architecture and crosstalk compensation method can also be implemented by global driving of entire cells using similar techniques to those in Refs. [12, 14].
VII Mid-circuit measurements
The ability to perform mid-circuit projective measurements, and apply coherent feedback based on the measured results, is at the heart of QEC, as well as additional central quantum computational tools [68, 69, 70]. In trapped-ion based systems, projective state detection is typically performed by state-dependent florescence. This poses a technical challenge, as the scattered photons usually heat-up the ion-crystal due to photon recoil. Furthermore the photon may be resonant with neighboring ions and scattered by them, resulting in decoherence of the quantum state of the entire system. Thus, so far, mid-circuit measurements have been implemented in trapped ions either in small ion-crystals, which are well separated from other parts of the quantum register, by ‘shelving’ all computational qubits to non-resonant atomic states, or by using two different atomic species spectraly separating logical and incoherent operations [71, 72, 73, 74].
In our optically segmented ion-crystal scheme, optically confined ions are well-suited for mid-circuit measurements, as their motion is separated to independent bands (top part of the spectra in Figs. 4 and 5), thus photon scattering heating remains local and does not heat the other crystal modes. Furthermore, the optical fields generating the trapping also substantially light-shift the atomic transition lines of the ions, specifically the photon emission lines, such that photon scattering is largely suppressed. Thus no physical shuttling of the ions into dedicated measurement regions is required, nor the use of multiple ion-species.
Figure 8: Top: schematic of mid-circuit measurement protocol, showing three types of involved ions, computational qubit ions (top, purple circle), barrier ion, $B_{\text{B}}$, to be measured (middle, purple circle) and optically confined barrier ions, $B_{\text{A}}$ (bottom, purple full circle). Operations are ordered from left to right. Data is encoded on the computational qubit (e.g. $U_{s,n}^{\left(A\right)}$), and one of the qubit states of the ion is ‘shelved’ to a non-fluorescing state using local control. Next an intermediate optical segmentation configuration is used, confining both ions A and B barrier ions, and separating their motion from the computational qubits. The state of the barrier ions is measured and then reset (‘Prepare’). Lastly the optical segmentation configuration is switched. Bottom: Relevant atomic levels of ${}^{40}\text{Ca}^{+}$ ions, showing $4S_{1/2}$ the ground state manifold, the metastable $4D_{5/2}$ manifold and the short-lived $4P_{1/2}$ manifold. The $S\leftrightarrow P$ transition varies between the no-optical confinement (left) and optically confined (right) cases, allowing to separate the two cases spectrally (see main text). The main optical field used in the protocol are shown, namely at $729$ nm (red), coupling the $S\leftrightarrow D$ levels, at $400$ nm (dark-blue), generating Raman transition between qubit states as well as additional local light-shifts, and at $397$ nm (light-blue) coupling the $S\leftrightarrow P$ levels.
We describe the protocol used to measure the qubit state of some target ions, without incurring decoherence on computational qubits, or heating of bulk modes. After the measurement is completed, we cool and re-initialize the measured ions. Figure 8 (top) illustrates these general steps involved in performing mid-circuit measurements. For simplicity we consider the required operations on three types of ions in the ion-crystal, i.e. computational qubit ions (top, purple circles), barrier ions to be measured (middle, purple circle), and optically confined barrier ions (bottom, purple full circles). Crucially, our measurement method can be performed in parallel throughout the entire ion-crystal on all required ions according to these roles. In our architecture ions may be optically confined specifically for the purpose of mid-circuit measurements, without having an essential role in segmentation of the ion-crystal, as showcased in the application of a QEC surface code, above.
For concreteness we consider a specific realization, in trapped ${}^{40}\text{Ca}^{+}$ ions. Furthermore we consider local control over a wavelength $\sim 400$ nm, which couples the two qubit states defined in the $4S_{1/2}$ ground state manifold via Raman transitions mediated by the short lived $4P$ manifolds. This local control, acting independently on all ions, is primarily used for generating single qubit rotations and multi-qubit programmable gates; however here it is utilized as a means to ‘localize’ global control fields using light-shifts, detailed below.
The main atomic levels used for standard qubit operation are presented in Fig. 8 (bottom) in the case of no optical confinement (left, red lines) and with optical confinement (right, blue lines). The ground state manifold is coupled to the narrow meta-stable $4D_{5/2}$ manifold via a $729$ nm field (red arrow). In the case of no optical confinement, this transition can be addressed on specific individual ions by light-shifting the $4S_{1/2}$ manifold of a target ion by using the locally controlled $400$ nm field (blue arrow). For these purposes a small, e.g. few MHz, light-shift suffices for individual ions addressing, while maintaining a low decoherence rate. On the other hand, optical confinement strongly light-shifts the $4S_{1/2}\leftrightarrow 4P_{1/2}$ level (right), making the $729$ nm coupling off resonant. In principle, this strong shift can be taken into account, however since the $S\leftrightarrow D$ transition is narrow, it imposes stringent requirements on the intensity stabilization of the optical confinement. Here we simply do not make use of the $S\leftrightarrow D$ transition for optically confined ions, relaxing the requirements for intensity stabilization. Raman transitions within the qubits subspace, mediated by the $400$ nm field, are negligibly affected by optical confinement, and exclusively in terms of their Rabi frequency.
An additional locally-applied field resonant with the light-shifted $S\leftrightarrow P$ transition at $397$ nm is used (purple arrow). Since the $P$ level is broad, the required stability of the optically confining fields in order to maintain resonance is reasonable. The combination of local $400$ nm and $397$ nm fields, together with standard repump fields (not shown) accommodates for all the necessary steps - i.e., state detection, various cooling techniques, and state preparation. We note that the technical requirements for the $397$ nm field are less demanding than those of the optical confinement and gate drive fields at $400$ nm; the necessary power per ion is far lower, independent spectral control is not required, and the focal spot can generally be larger (by illuminating adjacent ions within the same beam).
The mid-circuit measurement protocol shown in Fig. 8 is carried out by first encoding some information on ion B via the entanglement operations detailed above ( $U_{s,n}^{\left(A\right)}$), e.g. parity information of a plaquette shown in Fig. 3. Then, one of the qubit states of the ion to be measured is ‘shelved’ to the D manifold via a combination of $400$ nm light-shift and $729$ nm coupling, making this operation local on the ion.
Next, the segmentation configuration is changed to an intermediate setting where all barrier ions are illuminated. The purpose of optically confining the ion is to prepare it for being measured, by separating its motion from the computational qubits’ motion and to light-shift its $S\leftrightarrow P$ transition. We note that, for this protocol, the strength of the optical trapping potential should be adjusted so that all of the barrier bands are fully separated from the bulk bands (slightly above the value marked by the horizontal line in Fig. 4). Moreover, it is possible that the ion fully populates in the $D$ level, in which case it is not optically confined; thus existing confinement of other ions (purple full circle) cannot be relaxed. State detection is performed using the light-shifted $397$ nm field, followed by qubit reset and preparation using a combination of the $397$ nm and $400$ nm fields.
During the detection and preparation steps, the optically confined ions scatter $397$ nm photons, such that they are required to be cooled down in the ‘prepare’ step, which can be done in parallel. The scattered photons are substantially detuned, by approximately $0.8$ GHz, from the absorption lines of the computational qubit ions, having a linewidth of few tens of MHz, thus negligibly influencing their state. Nevertheless the state of neighboring computational qubits, located several micrometers apart, may be further protected by shelving both of their qubit states to the $D$ level (at the shelving step). Alternatively they may simply remain unutilized.
Lastly the segmentation configuration can now be changed, exchanging the roles of computational and barrier ions. The cooled ions are returned to the bulk of the cells and in turn cool down the bulk modes via sympathetic cooling.
VIII Summary
Here we propose a scalable architecture for quantum computing; it is based on a large register of trapped-ion qubits together with dynamically operated optical potentials. Our proposed architecture circumvents the two most prominent challenges in working with ever-larger ion-crystals – prohibitively high heating rates and spectral crowding of the ions’ motional modes. It does so by effectively segmenting an arbitrarily large trapped-ion crystal into several independent segments of a manageable size. Connectivity across the full trapped-ion crystal is enabled by rapidly reconfiguring the optical potentials. The optical potentials further enable mid-circuit measurements of the confined ions, followed by classical feedback.
The utility of this architecture is emphasized when combined with a method for programmable multi-qubit entangling gates, such as that proposed in Ref. [14]. We have used this method to numerically study the application of independent multi-qubit unitaries on each cell in parallel and in a scalable manner. Moreover we have extended this method to enable arbitrarily good compensation of crosstalk errors that arise between adjacent segments.
Our architecture requires modest hardware resources and makes use of well-established experimental techniques; it is thus an ideal platform for quantum simulation, near-term quantum computation, as well as ultimately for fault tolerant quantum computing.
Acknowledgements.
This work was supported by the Israel Science Foundation Quantum Science and Technology (Grants 2074/19, 1376/19 and 3457/21).